# Stagewise Reinforcement Learning and the Geometry of the Regret Landscape

### v1

@@ 5.3 Relevance to alignment
The results of Section 2 have implications for AI alignment. There is a subtle but important distinction between the prevailing (albeit implicit and underdeveloped) view on how policies of deep RL agents develop over training, and the SLT-informed view that this paper provides:
@@
In cases where there are any theoretical guarantees at all, alignment assumes “near enough” convergence to an optimal policy, which is safe (Hadfield-Menell et al., 2024; Kosoy, 2023). Let $\delta G$ be the allowed gap to this optimal safe policy and let us refer to any policy with higher regret as “misaligned”. We analogize Bayesian learning to SGD training, under some relation between dataset size $n$ and the number of trajectories seen over training $BS$ as in Section 5.2, and let $C=n/\log n$ be determined by the available compute. Then by (17) the generalized posterior can prefer a local minimum of $G$ which is misaligned but simpler than the optimal (aligned) policy by more than $C\delta G$ (Pepin Lehalleur et al., 2025).
Goal misgeneralization (Langosco et al., 2022; Abdel Sadek et al., 2025) is arguably an example of this phenomena. Indeed Langosco et al. (2022, §2.2) give inductive bias towards simplicity as their main suggestion for the origin of goal misgeneralization. Our experimental setting is a simplified form of the one studied in that work, and in the training run presented in Figure 1 if training had stopped at checkpoint 2k, the policy would be “misaligned” in the sense that the designer of the environment might have intended a trained agent to seek the cheese (this is after all the optimal policy).
Similar logic applies to reward hacking in RLHF (Lambert, 2026). Reward models trained on human preference data may learn simple correlations (e.g., “longer responses are better” (Shen et al., 2023) or “lists are preferred” (Zhang et al., 2025)) rather than complex representations of genuine quality (Gao et al., 2022). If the simpler pattern achieves comparable loss on the preference dataset, the free energy formula (for supervised learning) predicts it will be favored during training. When this reward model is used to train the policy, these biases are amplified and also interact with the simplicity bias of the RL training process.
@@ 6 Conclusion
Our empirical results validate the theory’s predictions. In a gridworld environment exhibiting stagewise learning, we observe “opposing staircases” where phase transitions manifest as rapid decreases in regret accompanied by rapid increases in the LLC. Notably, the LLC detects the transition between phases 2b and 3 even when estimated on states where both policies achieve identical regret, suggesting it captures changes in the underlying algorithm rather than merely performance. This demonstrates that the geometry of the regret function reveals structure in the learning process that is invisible to behavioral metrics alone.
@@
Acknowledgements
We would like to thank Edmund Lau for his involvement in the early stages of this project and for many helpful suggestions. We are very grateful to Marcus Hutter and Vanessa Kosoy for productive conversations on the topic of this paper during the Mathematical Science of AI Safety focus period at SMRI. We are also grateful to Rohan Hitchcock for sharing the argument used in the proof of Lemma E.31 and to Shaowei Lin for providing guidance concerning the control as influence literature. Finally we would like to thank Zach Furman, Philipp Alexander Kreer and Rumi Salazar for their feedback on an earlier draft.
@@ Acknowledgements
Appendix A Outlook
In this appendix we give a discussion of some further directions that the introduction of singular learning theory to the reinforcement learning context allows one to pursue.
Weight-Refined LLCs
It was demonstrated in Wang et al. [2025] that one can detect the development of structure in parameter space during training by studying what they refer to as weight-refined LLCs – LLCs for the restriction of the loss function to a subspace of parameter space. They study transformer models and detect differentiation of attention heads during training using this method, and these techniques are also applicable in the reinforcement learning context. By estimating refined LLCs for, for instance, individual layers of the deep neural network in deep reinforcement learning we may detect development of internal structure that is not visible from looking at the policy alone.
Susceptibilities
The techniques of singular learning theory may be applied to the expectation values of more general observables with respect to the Gibbs posterior. What’s more, if the loss function one considers depends on a hyperparameter (in our setting we saw the dependence of the regret on the hyperparameters $\alpha,\gamma$) one can study the infinitesimal change in these expectation values with respect to the variation of these hyperparameters. In Baker et al. [2025] the authors study exactly this setting: their susceptibilities correspond to the variation of weight-refined LLCs under infinitesimal perturbations of the data distribution. In the reinforcement learning setting one can do the same thing, where one varies either the reward function of the Markov problem or the initial state distribution.
Environment Complexity
By enhancing our toy model slightly one could explore the ability of the LLC estimator to detect out-of-distribution generalization, meaning here performance in classes of states not seen during training. For a simple example consider the extension of the model presented here to maze environments (as in the original cheese-in-the-corner example presented in Abdel Sadek et al. [2025]). One can allow the distribution of initial states used during training and LLC estimation to differ, as we did in Section 3.4.4, but now changes in the LLC may detect changes in performance in out-of-distribution states.
Stochastic Transitions
Finally we discuss a theoretical question. To what extent is it possible to relax the assumption that the transition functions of the Markov problem are deterministic? We may immediately relax it a little to the condition given in Assumption E.16 – that optimal policies may obtain optimal reward with probability one – but new techniques will be required to go beyond this. Note that this condition is analogous to a hierarchy of assumptions that appear in Watanabe’s work in the context of distribution learning from iid data (essential uniqueness, renormalizability as in Watanabe [2010] or the relative finite variance condition in Watanabe [2018]). An approach to generalize beyond this might build on the non-essentially-unique example of Nagayasu and Watanbe [2022], in which a more significant term appears in the free energy.
Appendix B LLC Estimation
@@
Appendix C Predicting transitions
@@
Now, all the values of $\gamma\in\{0.9,0.950.975,0.99\}$ are relatively close, but using the value of $\gamma$ directly as a feature doesn’t reflect the hyperbolic nature of the discount factor, as $\gamma\to 1$, the agent becomes “infinitely” far-sighted. So, we instead compute the effective horizon $h=\frac{1}{1-\gamma}$, and model the number of gradient steps taken for the transition $\pi^{1}\to\pi^{2b}$ linearly as
@@
for learned coefficients $c_{1},c_{2},c_{3}$. We find the best fit to be $175.52+459.20\alpha+12.00h$, with a coefficient of determination of $R^{2}=0.870$, and a relative error of at worst $\pm 30\%$.
@@
Appendix D Model Architecture
@@
The model follows an IMPALA-style convolutional encoder with residual blocks presented in Abdel Sadek et al. [2025], see Figure 8. The previous action is embedded as a one-hot vector of length 4 (with the zero vector encoding the lack of a previous action on the first timestep). This then is fed into a small feedforward network (in place of a LSTM) which then feeds into a policy head, from which we obtain the logits over the actions.
The original network in Abdel Sadek et al. [2025] also adds a value head, which has been removed as we make no use of it. At a high level the network maps an input observation $x\in\mathbb{R}^{C\times H\times W}$ and a previous action $a_{t-1}\in\{\uparrow,\downarrow,\leftarrow,\rightarrow,\varnothing\}$ to policy logits $\pi(\cdot|s_{t},a_{t-1})\in\mathbb{R}^{4}$, which can be softmaxed to obtain a probability distribution over actions.
D.1 Convolutional trunk
@@
D.2 Previous-action Embedding
The previous action is encoded as a $4$-dimensional one-hot vector; the null previous action $a_{0}=\varnothing$ is encoded as the zero vector. This action embedding is concatenated with the observation embedding output from the encoder stack, and then fed into the last feature extractor layer $\mathrm{Linear}(260\to 256)$.
D.3 Heads
@@
D.4 Hyperparameters
@@ D.5 Compute
Appendix E Theoretical Foundations of Singular Learning Theory in Reinforcement Learning
E.1 Markov Problems and the Regret
E.1.1 Initial Setup
@@
Remark E.2.
In Toussaint and Storkey [2006] Toussaint and Storkey study Markov decision problems of unbounded length by considering an infinite mixture of bounded length decision problems. We will not pursue this approach in the formalization considered below but we expect that the techniques will generalize to their setting.
Definition E.3.
@@
E.1.2 Importance Weighting Estimators
@@
Observe that Assumption E.8 guarantees that the random variables $G_{n}(w)$ are real analytic function valued. That is, $G_{n}$ is real analytic as a function
@@
E.2 Singular Learning Theory for Generalized Bayesian Inference
@@
E.2.1 Setup for Generalized Bayesian Inference
In this section we will describe a general setting for generalized Bayesian inference that allows us to study singular learning theory for supervised learning and reinforcement learning, among other situations, in a unified fashion. We will see shortly in Example E.20 that the Markov decision problem setting introduced in the previous section is an example of the situation we will describe now. Our basic setting will comprise the following.
@@ E.2.1 Setup for Generalized Bayesian Inference
Example E.19.
Our first example is considered in Watanabe’s original work Watanabe [2009]: supervised learning for a model where the true distribution is realizable. Choose any compact parameter space $W$ and sample space $X$ and specify a real analytic family $p_{w}$ of probability distributions on $X$.
@@
that is, the loss agrees with the Kullback–Leibler divergence from the “true distribution” $p_{w^{*}}$. A point $w$ is optimal if $p_{w}=p_{w^{*}}$. If so then $g(x)q_{w}(x)-h(x)=0$ for all $x$, so Assumption E.16 holds automatically.
For the estimators in this example we generalize Definition E.18 slightly in this case and instead consider $x_{i}$ samples from $p_{w^{*}}$ for $i=1,\ldots n$ with
@@
Example E.20.
Now let us consider a Markov decision problem as in E.1 and let $X=\mathcal{T}$ be the space of trajectories. Let $\Pi$ be the space of policies, and choose an analytic parameterization map $F\colon W\to\Pi$ sending $w$ to $\pi_{w}$. Let $q_{w}$ be the distribution on $\mathcal{T}$ associated to this choice of policy. If $\tau\in\mathcal{T}$, let $g(\tau)$ be defined as in Definition E.10 and let $h(\tau)=0$.
Assumption E.16 is not necessarily satisfied; it holds if optimal policies pursue optimal trajectories with probability one. We will discuss this further in Section E.2.4 below.
Note that in the Markov decision problem setting Assumption E.16 holds in particular if the decision problem has deterministic transition functions, meaning that the distributions $p(o|s)$ and $p(s|a)$ are all delta distributions. Say a trajectory $\tau$ is reachable for an MDP if there exists $w\in W$ so that $q(w,\tau)>0$.
Proposition E.21.
Let $\pi_{w}$ be an optimal policy in a Markov decision problem. If the problem has deterministic transition functions then $g(\tau)\geq 0$ for all reachable trajectories $\tau\in\mathcal{T}$ and $q_{w}(\tau)=0$ if and only if $g(\tau)>0$. In particular Assumption E.16 holds.
@@
Remark E.23.
The generalized posterior with $\beta=1$ is determined canonically by Theorem 1 of Bissiri, Holmes and Walker Bissiri et al. [2016]. Indeed, given a prior distribution $\phi(w){\rm d}w$ and the loss function $f(w,w^{\prime},x)$ we consider the collection of all probability measures of the form
@@
Their proof applies in the case where $W$ is countable and discrete, but we note that since the function $\psi(f(w,w^{\prime},x_{2}),\psi(f(w,w^{\prime},x_{2}),\phi(w))))-\psi(f(w,w^{\prime},x_{1})+f(w,w^{\prime},x_{2}),\phi(w))$ is continuous in $w\in W$ it is enough to show that it vanishes on a countable dense subset of $W$. Together with Remark E.26 below we see that the generalized tempered posterior is canonically characterized by the loss function under natural assumptions.
@@
E.2.2 Convergence of Error Terms
As the first step in our proof of Theorem E.24 we will need to establish a version of the results from [Watanabe, 2009, Section 5] that applies in the generalized Bayesian inference setting in which our samples $x_{i}$ are not identically distributed. We will first establish that a version of the central limit theorem my be applied to our importance weighted regret estimators.
Lemma E.25.

### v2

@@ 7.1 Relevance to Alignment
Our results have implications for AI alignment. There is a subtle but important distinction between the prevailing (albeit implicit and underdeveloped) view on how policies of deep RL agents develop over training, and the SLT-informed view that this paper provides:
@@
In cases where there are any theoretical guarantees at all, alignment assumes “near enough” convergence to an optimal policy, which is safe (Hadfield-Menell et al., 2024; Kosoy, 2023). Let $\delta G$ be the allowed gap to this optimal safe policy and let us refer to any policy with higher regret as “misaligned”. We analogize Bayesian learning to SGD training, under some relation between dataset size $n$ and the number of trajectories seen over training $B\cdot S$ as in Section 5.4, and let $C=n/\log n$ be determined by the available compute. Then by (17) the generalized posterior can prefer a local minimum of $G$ which is misaligned but simpler than the optimal (aligned) policy by more than $C\delta G$ (Pepin Lehalleur et al., 2025).
Goal misgeneralization (Langosco et al., 2022; Shah et al., 2022; Abdel Sadek et al., 2025) is arguably an example of this phenomena. Indeed Langosco et al. (2022, §2.2) give inductive bias towards simplicity as their main suggestion for the origin of goal misgeneralization. Our experimental setting is a simplified form of the one studied in that work, and in the training run presented in Figure 1 if training had stopped at checkpoint 2k, the policy would be “misaligned” in the sense that the designer of the environment might have intended a trained agent to seek the cheese (this is after all the optimal policy).
Similar logic applies to reward hacking in RLHF (Lambert, 2026). Reward models trained on human preference data may learn simple correlations (e.g., “longer responses are better” Shen et al., 2023 or “lists are preferred” Zhang et al., 2025) rather than complex representations of genuine quality (Gao et al., 2022). If the simpler pattern achieves comparable loss on the preference dataset, the free energy formula (for supervised learning) predicts it will be favored during training. When this reward model is used to train the policy, these biases are amplified and also interact with the simplicity bias of the RL training process.
@@
7.2 Model Complexity in Reinforcement Learning
In the (generalized) Bayesian setting, SLT tells us that the LLC provides a rigorous formalization of the notion of complexity of models. Our empirical results are evidence that this notion of complexity is also applicable in the context of stochastic optimization and deep RL. In this section, we survey related work on model complexity and reinforcement learning.
Measurements of complexity, including those quantifying flatness of a local minimum in the loss landscape have been widely studied for neural networks (Hochreiter and Schmidhuber, 1997; Keskar et al., 2016; Neyshabur et al., 2017; Dherin et al., 2022). Likewise many authors have explored the idea of simplicity bias in neural networks as an implicit regularization affecting generalization (Shah et al., 2020; Teney et al., 2025; Pérez et al., 2018; Morwani et al., 2023). The notion of the loss-complexity tradeoffs and transitions between different solution classes is studied in Singh et al. (2023); Munn and Wei (2025); Carroll et al. (2025); Wurgaft et al. (2025).
The direct study of complexity bias in reinforcement learning specifically has been less explored. The geometry of the reward landscape was explored visually in many examples in Sullivan et al. (2022), which demonstrated the appearance in low-dimensional cross-sections of “cliffs” and “plateaus” in the landscape. Boucher et al. (2024) studied the inductive bias introduced by the technique of maximum entropy learning (Ziebart et al., 2008) using a range of complexity measures including those that quantified local flatness of a minimum. This idea of flatness specifically as a bias favoring more robust optima was investigated in Lee and Yoon (2025). Boopathy et al. (2023) studied inductive biases in a range of contexts including reinforcement learning using an information-theoretic measure of complexity. Other authors have applied complexity bias to improve learning performance in reinforcement learning problems (You et al., 2025; Lee et al., 2024), while others have studied the generalization of RL agents in grid world similar to the environments considered in this paper (Ramanauskas and Şimşek, 2023; Abdel Sadek et al., 2025).
We lastly mention some of the theoretical literature on the dynamics of reinforcement learning. The dynamics under policy gradient learning was studied in an illustrative model for reinforcement learning in Patel et al. (2025). The connection between complexity and stagewise development has recently been explored in the supervised learning case, particularly for large language models (Wei et al., 2022a; Chen et al., 2024; Wang et al., 2024; Hoogland et al., 2025).
7.3 Open Questions and Opportunities
In this section, we give a discussion of some further directions that the extension of singular learning theory to the reinforcement learning context allows one to pursue.
7.3.1 Dynamics of Deep Reinforcement Learning
Understanding the inductive biases of deep RL agents is a fundamental challenge for the science of deep learning.
In this paper, we have bridged the first part of the gap in our current understanding. By extending singular learning theory to an RL setting, we offer the first theory that accounts for the role of geometric degeneracy in reinforcement learning. The qualitative correspondence between stagewise development in our theoretical model and our experimental results is evidence that our theoretical framework has captured something of the fundamental role played by degeneracy in SGD-based deep RL.
This is progress in the right direction, but, as we have discussed, it remains an open question to develop a complete theoretical account of the inductive biases of deep RL proper. In particular, we still need to account for the role played by a local, gradient-based, stochastic optimisation, as compared to the idealized learning process of progressive generalized posterior updates. Our results suggest that a promising path forward lies in studying the geometry of degenerate critical points, such as with the algebraic geometer’s toolkit.
7.3.2 Fundamental Conditions and Stochastic Transitions
Another important theoretical question is the following. To what extent is it possible to relax the assumption that the transition functions of the Markov problem are deterministic?
We may immediately relax this condition a little to the condition given in D.16 – that optimal policies may obtain optimal reward with probability one – but new techniques will be required to go beyond this. Note that this condition is analogous to a hierarchy of assumptions that appear in Watanabe’s work in the context of distribution learning from iid data (essential uniqueness, renormalizability as in Watanabe, 2010 or the relative finite variance condition in Watanabe, 2018). An approach to generalize beyond this might build on the non-essentially-unique example of Nagayasu and Watanbe (2022), in which a more significant term appears in the free energy.
7.3.3 Interpretability and Patterning via Weight-Refined LLCs and Susceptibilities
The extension of SLT to the RL setting also opens the door to some promising practical research directions, based on recent progress applying SLT to develop interpretability and training tools in the supervised learning setting.
Wang et al. (2025) demonstrated that one can detect the development of structure in parameter space during training by studying so-called weight-refined LLCs – LLCs for the restriction of the loss function to a subspace of parameter space. They study transformer models and detect differentiation of attention head roles during training using this method. Our work shows that these techniques are also applicable in the reinforcement learning context. By estimating refined LLCs for, for instance, individual layers of the deep neural network in deep reinforcement learning we may detect development of internal structure that is not visible from looking at the policy alone.
The techniques of singular learning theory may be applied to the expectation values of more general observables with respect to the Gibbs posterior. What’s more, if the loss function one considers depends on a hyperparameter (in our setting we saw the dependence of the regret on the hyperparameters $\alpha,\gamma$) one can study the infinitesimal change in these expectation values with respect to the variation of these hyperparameters. Baker et al. (2025) study exactly this setting for supervised learning: their susceptibilities correspond to the variation of weight-refined LLCs under infinitesimal perturbations of the data distribution. In the reinforcement learning setting one can do the same thing, where one varies either the reward function of the Markov problem or the initial state distribution.
One can take this a step further. Wang and Murfet (2026) introduce a technique, patterning, by which one may use susceptibilities to steer the training process in an optimal direction for the promotion of desired behaviors, as measured by the maximization or minimization of the posterior expectation value of specified observable quantities. In other words, one may take advantage of the ability of susceptibilities to measure the dependence of the local geometry of the loss landscape on parameters – such as the mixing parameter in a mixture of data distributions – to modify this geometry in order to promote desirable optima. In the reinforcement learning context one can use the results we have developed in order to apply the same techniques to modify the favorability of different regions in the space of optima by perturbing structures of the Markov problem such as the initial state distribution. The result would be a method of automatically adapting the RL agent’s training distribution so as to promote certain desirable kinds of agent behavior, comparable to unsupervised environment design, as used to elicit robust capability- or goal-generalization (cf., Dennis et al., 2020; Abdel Sadek et al., 2025).
@@ 8 Conclusion
Our empirical results validate the theory’s predictions. In a gridworld environment exhibiting stagewise learning, we observe “opposing staircases” where phase transitions manifest as rapid decreases in regret accompanied by rapid increases in the LLC.
@@
Acknowledgments
We would like to thank Edmund Lau for his involvement in the early stages of this project and for many helpful suggestions. We are very grateful to Marcus Hutter and Vanessa Kosoy for productive conversations on the topic of this paper during the Mathematical Science of AI Safety focus period at SMRI. We are also grateful to Rohan Hitchcock for sharing the argument used in the proof of Lemma D.31 and to Shaowei Lin for providing guidance concerning the control as influence literature. Finally we would like to thank Zach Furman, Philipp Alexander Kreer and Rumi Salazar for their feedback on an earlier draft.
@@ Acknowledgments
Appendix A LLC Estimation
@@
Appendix B Predicting transitions
@@
Now, all the values of $\gamma\in\{0.97,0.98,0.99\}$ are relatively close, but using the value of $\gamma$ directly as a feature doesn’t reflect the hyperbolic nature of the discount factor, as $\gamma\to 1$, the agent becomes “infinitely” far-sighted. So, we instead compute the effective horizon $h=\frac{1}{1-\gamma}$, and model the number of gradient steps taken for the transition $\pi^{1}\to\pi^{2b}$ linearly as
@@
for learned coefficients $c_{1},c_{2},c_{3}$. We find (Figure 5) the best fit to be $180.2+687.9\alpha+15.1h$, with a coefficient of determination of $R^{2}=0.993$, and a relative error of at worst $\pm 9.1\%$.
@@
Appendix C Model Architecture
@@
The model follows an IMPALA-style convolutional encoder with residual blocks presented in Abdel Sadek et al. [2025], see Figure 6. This then is fed into a small feedforward network (in place of a LSTM) which then feeds into a policy head, from which we obtain the logits over the actions.
The original network in Abdel Sadek et al. [2025] also adds a value head, which has been removed as we make no use of it. At a high level the network maps an input observation $x\in\mathbb{R}^{C\times H\times W}$ to policy logits $\pi(\cdot|s_{t})\in\mathbb{R}^{4}$, which can be softmaxed to obtain a probability distribution over actions.
C.1 Convolutional trunk
@@
C.2 Heads
@@
C.3 Hyperparameters
We provide our hyperparameter selections in Table 1.
@@ C.4 Compute
Appendix D Theoretical Foundations of Singular Learning Theory in Reinforcement Learning
In this appendix, we prove the results stated in Section 4. Section D.1 contains preliminary definitions, including a formulation of analytic Markov decision processes that relaxes the finiteness conditions adopted in the main text for brevity. Section D.2 introduces a unified definition of an inference problem that subsumes both the generalized Bayesian inference for RL developed in the main text and also the supervised learning setting of Watanabe, among other settings, and then proves the main results in this unified setting.
D.1 Markov Problems and the Regret
D.1.1 Initial Setup
@@
Remark D.2.
Toussaint and Storkey [2006] study Markov decision problems of unbounded length by considering an infinite mixture of bounded length decision problems. We will not pursue this approach in the formalization considered below but we expect that the techniques will generalize to their setting.
Definition D.3.
@@
D.1.2 Importance Weighting Estimators
@@
Observe that Assumption D.8 guarantees that the random variables $G_{n}(w)$ are real analytic function valued. That is, $G_{n}$ is real analytic as a function
@@
D.2 Singular Learning Theory for Generalized Bayesian Inference
@@
D.2.1 Setup for Generalized Bayesian Inference
In this section we will describe a general setting for generalized Bayesian inference that allows us to study singular learning theory for supervised learning and reinforcement learning, among other situations, in a unified fashion. We will see shortly in Example D.20 that the Markov decision problem setting introduced in the previous section is an example of the situation we will describe now. Our basic setting will comprise the following.
@@ D.2.1 Setup for Generalized Bayesian Inference
Example D.19.
Our first example is considered in Watanabe’s original work [Watanabe, 2009]: supervised learning for a model where the true distribution is realizable. Choose any compact parameter space $W$ and sample space $X$ and specify a real analytic family $p_{w}$ of probability distributions on $X$.
@@
that is, the loss agrees with the Kullback–Leibler divergence from the “true distribution” $p_{w^{*}}$. A point $w$ is optimal if $p_{w}=p_{w^{*}}$. If so then $g(x)q_{w}(x)-h(x)=0$ for all $x$, so Assumption D.16 holds automatically.
For the estimators in this example we generalize Definition D.18 slightly in this case and instead consider $x_{i}$ samples from $p_{w^{*}}$ for $i=1,\ldots n$ with
@@
Example D.20.
Now let us consider a Markov decision problem as in D.1 and let $X=\mathcal{T}$ be the space of trajectories. Let $\Pi$ be the space of policies, and choose an analytic parameterization map $F\colon W\to\Pi$ sending $w$ to $\pi_{w}$. Let $q_{w}$ be the distribution on $\mathcal{T}$ associated to this choice of policy. If $\tau\in\mathcal{T}$, let $g(\tau)$ be defined as in Definition D.10 and let $h(\tau)=0$.
Assumption D.16 is not necessarily satisfied; it holds if optimal policies pursue optimal trajectories with probability one. We will discuss this further in Section D.2.4 below.
Note that in the Markov decision problem setting Assumption D.16 holds in particular if the decision problem has deterministic transition functions, meaning that the distributions $p(o|s)$ and $p(s|a)$ are all delta distributions. Say a trajectory $\tau$ is reachable for an MDP if there exists $w\in W$ so that $q(w,\tau)>0$.
Proposition D.21.
Let $\pi_{w}$ be an optimal policy in a Markov decision problem. If the problem has deterministic transition functions then $g(\tau)\geq 0$ for all reachable trajectories $\tau\in\mathcal{T}$ and $q_{w}(\tau)=0$ if and only if $g(\tau)>0$. In particular Assumption D.16 holds.
@@
Remark D.23.
The generalized posterior with $\beta=1$ is determined canonically by Theorem 1 of Bissiri et al. [2016]. Indeed, given a prior distribution $\phi(w){\rm d}w$ and the loss function $f(w,w^{\prime},x)$ we consider the collection of all probability measures of the form
@@
Their proof applies in the case where $W$ is countable and discrete, but we note that since the function $\psi(f(w,w^{\prime},x_{2}),\psi(f(w,w^{\prime},x_{2}),\phi(w))))-\psi(f(w,w^{\prime},x_{1})+f(w,w^{\prime},x_{2}),\phi(w))$ is continuous in $w\in W$ it is enough to show that it vanishes on a countable dense subset of $W$. Together with Remark D.26 below we see that the generalized tempered posterior is canonically characterized by the loss function under natural assumptions.
@@
D.2.2 Convergence of Error Terms
As the first step in our proof of Theorem D.24 we will need to establish a version of the results from Watanabe [2009, §5] that applies in the generalized Bayesian inference setting in which our samples $x_{i}$ are not identically distributed. We will first establish that a version of the central limit theorem my be applied to our importance weighted regret estimators.
Lemma D.25.